Rita has a collection of 108 coins consisting of nickels, dimes, and quarters. The number of dimes is 8 more than one-third the number of nickels, and the number of quarters is twice the number of dimes. How many coins of each kind does she have?
Please help me solve this

Answer by FrankM(1040)

N+D+Q=108
D=(N/3)+8
Q=2D
Since Q=2D, D+Q= 3D so go to D=(N/3)+8 and triple it.
3D=N+24
and
N+N+24 = 108
2N= 84
N=42
Since 3D = N+24
3D= 66, and D=22
Q= 108-42-22= 44
N=42, D=22, Q=44
Check and see the numbers solve for each of the original criteria.